# Semi-Supervised Unconstrained Head Pose Estimation in the Wild

### v3

@@ Abstract
publicationid: pubid: $0$–$00$ 2025 IEEE
@@ 1 Introduction
In the era of deep supervised learning, many HPE methods [17, 18, 19, 20] need a large amount of labeled data to train. However, existing HPE datasets such as 300W-LP [11], BIWI [21] and DAD-3DHeads [16] have their incompatible limitations for real applications. They are either artificially collected or synthesized [11, 21, 22, 23] so that having huge domain gaps and scarce diversities compared to natural heads. Others are manually annotated by certified experts [16, 24] at a significant time and economic cost with a small scale. Some examples are shown in Fig. 2. As a result, although many supervised learning methods designed and trained based on these datasets have achieved excellent quantitative performance on the test set (such as AFLW2000 [11] and BIWI [21]), these models cannot be directly applied to the real world with complex scenarios. Some methods try to synthesize a large number of multi-view [25], extreme-view [26] or full-view [27] human heads to expand the training set, but these generated images still have hallucination defects and cannot be fully trusted. Different from most existing fully supervised methods, we turn to the semi-supervised learning (SSL) techniques [28, 29, 30, 31], and propose a semi-supervised unconstrained head pose estimation (SemiUHPE) method, that can leverage extensive easier obtainable yet unlabeled in-the-wild heads [32, 33, 34, 35] in addition to partially labeled data to improve performance and generalization. The overall framework of our method is illustrated in Fig. 3. It can not only avoid the laborious annotation of 3D head pose on 2D images which itself is ill-posed, but also greatly promote the estimation accuracy of challenging cases in real environments.
@@
In method section, we provide more details of above-mentioned strategies. We also present how these strategies can be utilized to solve the task generic object rotation regression similar to HPE and help improve another task 3D head reconstruction strongly related to HPE. Despite the simplicity, these seamless adaptations reveal valuable universality and scalability of SemiUHPE. In experiments, we demonstrate the effectiveness and versatility of our proposed strategies by extensive comparison and ablation studies. In addition, we give impressive qualitative estimation results on challenging wild heads, which is promising for downstream applications.
@@ 2.3 Semi-Supervised Rotation Regression
This is a less-studied field compared with other popular SSL tasks such as classification and detection. For the 6D pose estimation, Self6D [69, 70] establishes self-supervision for 6D pose by enforcing visual and geometric consistencies on top of unlabeled RGB-D images. NVSM [71] builds a category-level 3D cuboid mesh for estimating pose of rigid object in a synthesis-and-matching way. Recently, based on FixMatch [30], FisherMatch [36] firstly proposes the semi-supervised rotation regression for generic objects. It can automatically learn uncertainties along with predictions by introducing the matrix Fisher distribution [38, 39] to build a probabilistic model of rotation. The entropy of this distribution has been validated to be an efficient indicator of prediction for pseudo label filtering. The improved version [37] has integrated the rotation Laplace distribution [72] which is more robust to the disturbance of outliers and enforces much gradient to the low-error region. Essentially, HPE is subordinate to the rotation regression problem. MFDNet [56] also utilizes the matrix Fisher distribution to model head rotation uncertainty. We follow [36, 56] and adopt this distribution to tackle the SemiUHPE task.
@@ 3.1 Framework Overview
$p(\mathbf{R})=\mathcal{MF}(\mathbf{R};\mathbf{A})=\frac{1}{F(\mathbf{A})}\exp(\text{tr}(\mathbf{A}^{T}\mathbf{R}))$ (1)
@@
$\mathbf{R}=\mathbf{U}\left[\begin{array}[]{ccc}1&\quad 0&\quad 0\\ 0&\quad 1&\quad 0\\ 0&\quad 0&\quad\det(\mathbf{U}\mathbf{V})\end{array}\right]\mathbf{V}^{T}$ (2)
@@
$H(f)=\log{F_{f}}-\sum\nolimits^{4}_{i=1}\left({z_{f_{i}}\frac{1}{F_{f}}\frac{\partial{F_{f}}}{\partial{z_{f_{i}}}}}\right)$ (3)
@@ 3.3 Dynamic Entropy-based Filtering
$\mathcal{L}_{unsup}(\mathbf{x}^{u})=\mathds{1}_{(H(p_{tea})\leq\tau)}\mathcal{L}^{CE}(p_{tea},p_{stu})$ (4)
@@
$\tau_{k}=\mathsf{percentile}\langle H(\mathcal{MF}(\mathcal{N}_{tea}^{k}(\mathbf{x}^{u}_{i})))|_{i=1}^{N_{u}},\delta\rangle$ (5)
@@
$\mathcal{L^{\prime}}_{unsup}(\mathbf{x}^{u})=\mathds{1}_{(H(p^{k}_{tea})\leq\tau_{k})}\mathcal{L}^{CE}(p^{k}_{tea},p_{stu})$ (6)
where $p^{k}_{tea}\!=\!\mathcal{MF}(\mathcal{N}_{tea}^{k}(\mathbf{x}^{u}))$. Usually, with a given percentage $\delta$, the dynamic entropy threshold $\tau_{k}$ will decrease as the stage $k$ increases. And the optimal value of $\delta$ is inversely proportional to the images number of $\mathcal{D}^{u}_{ood}$ in $\mathcal{D}^{u}$.
@@ 3.4.2 Pose-Altering Rotation Consistency
$\overline{p^{i}_{stu}}=\mathcal{MF}(\mathcal{N}_{stu}(\overline{\mathbf{x}^{u}_{i}})),\quad\overline{\mathbf{x}^{u}_{i}}=T_{\mathsf{CutOcc}}(T_{\mathsf{Rot}_{\theta}}(\mathbf{x}^{u}_{i}))$ (7)
$\widetilde{p^{i}_{tea}}=\mathcal{MF}(\mathcal{N}_{tea}(\widetilde{\mathbf{x}^{u}_{i}})),\quad\widetilde{\mathbf{x}^{u}_{i}}=T_{\mathsf{weak}}(\mathbf{x}^{u}_{i})$
$\widehat{p^{i}_{tea}}=\mathbf{M_{\theta}}\widetilde{p^{i}_{tea}},\quad\mathbf{M_{\theta}}=\small{\left[\begin{array}[]{ccc}\cos(\theta)&\sin(\theta)&0\\ -\sin(\theta)&\cos(\theta)&0\\ 0&0&1\end{array}\right]}$
@@ 3.5 Adaptation of SemiUHPE
Figure 8: The illustration of adapted SemiUHPE for solving the 3D head reconstruction task. The selected basic network is DAD-3DNet [16] with multiple outputs. The input part of two networks is omitted to save space. We additionally design a novel pose-guided filtering strategy for rejecting less-confident unlabeled samples.
@@
Now, we analyze how to filter pseudo labeled heads. After introducing a new branch, we can construct an empirically reasonable pose-guided pseudo-label filtering strategy. As shown in Fig. 9 left, the sample uncertainty can be reflected through the teacher model by calculating the difference between the sub-branch predicted rotation matrix and the 3DMM included rotation matrix (e.g., the Frobenius norm $\|\mathbf{I}\!-\!\mathbf{R}_{1}\mathbf{R}_{2}^{T}\|_{F}$). Alternatively, considering that computing matrix differences as an uncertainty indicator is not always statistically reliable, we modify the output of HPE sub-branch to construct a probabilistic rotation regression with matrix Fisher distribution (refer Fig. 9 right). There are two advantages to applying the matrix Fisher distribution. On one hand, its output is an arbitrary $3\!\times\!3$ matrix with homeomorphic parameters to $\mathbb{R}^{9}$ [39], which is easier to estimate than the 3D rotations in $\mathcal{SO}(3)$ topology [56]. On the other hand, we can measure the confidence of predicted distribution by calculating its entropy, and devise a self-reliant more robust filtering mechanism. We will verify that the adapted SemiUHPE can benefit from both filtering strategies, with the latter being superior.
@@ 3.6 Training Protocol
where $-\log(\cdot)$ means the negative log likelihood (NLL) of the mode predicted by $\mathcal{N}_{stu}$ in the distribution of ground-truth label $\mathbf{y}^{l}$. We save the best performed student model for cloning an identical teacher model for the next phase.
@@ 4.1 Datasets and Settings
Setting1: 300W-Self In this setting, we use partially annotated 300W-LP as the labeled set and the left part as the unlabeled set. The labeled ratio is selected from (2%, 5%, 10%, 20%). The Phase1 and Phase2 has 20K and 40K iterations, respectively. The number of threshold updating stages $K$ in Phase2 is 4. Batch size for the labeled set $B_{l}$ and unlabeled set $B_{u}$ is 32 and 128, respectively. The remained percentage $\delta$ of unlabeled data is 0.95. The test-set is AFLW2000. For a fair comparison, we keep samples with Euler angles within ±$\pm 90^{\circ}$ following the previous front-range HPE methods [18, 19, 20]. The evaluation metric is Mean Absolute Error (MAE) of Euler angles. We thus convert predictions into Euler angles for comparing.
@@
Compared Methods: In order to make a more comprehensive comparison, apart from our proposed SemiUHPE, we also implemented the fully supervised version (Sup.) using only labeled data but adopting an advanced matrix Fisher representation [39], and the semi-supervised baseline FisherMatch [36] (Base.) following our settings. For the network backbone, it is selected from three purely ConvNets-based candidates including ResNet50 [85], RepVGG [86] and EfficientNetV2-S [87]. These backbones have similar network parameters and are widely used in other HPE methods [18, 19, 20]. For all experiments, we implemented them using PyTorch 1.13.1 on one single RTX 3090 (24 GB) or A800 (80 GB) GPU. The Adam optimizer is used. The learning rate in Phase1 is 1e-4, and reduced to 1e-5 in Phase2. During training and testing, all head images are padded and resized into shape $224\times 224$ as inputs.
@@ 4.2 Implementation Details
$T_{\mathsf{CutOcc}}$ CutOut: 3 Holes CutMix: 3 Holes 1.0
@@ 4.3 Quantitative Comparison
Below, we report results under three settings as well as improvement analysis of our SemiUHPE. If there is no special reminder, both Sup. and Base. methods are implemented with our proposed aspect-ratio invariant cropping. The last is the results of SemiObjRot and Semi3DHead.
TABLE II: Euler angles errors on AFLW2000. Models are trained on 300W-LP with different ratios of label. The best result is in red color.
  Method Backbone 2% 5% 10% 20% All
  Sup. ResNet50 4.347 3.987 3.831 3.619 3.578
@@
[…]
  Method Reference Extra Pitch Yaw Roll MAE
  3DDFA [11] CVPR’16 3DMM 5.98 4.33 4.30 4.87
@@
Base. (ResNet50) CVPR’22 No 4.61 2.99 3.00 3.53
@@
  Method $\|\mathbf{I}\!-\!\mathbf{R}_{1}\mathbf{R}_{2}^{T}\|_{F}\!\downarrow$ Angle error (degree)$\downarrow$
@@
  Method $\|\mathbf{I}\!-\!\mathbf{R}_{1}\mathbf{R}_{2}^{T}\|_{F}\!\downarrow$   /   Angle error (degree)$\downarrow$
@@
  DAD-3DNet [16] 7.861 4.766 4.025 2.467 0.919 0.929 0.938 0.954 3.957 3.468 3.247 2.982 0.327 0.270 0.233 0.149
Mean-Teacher [28] 5.599 4.969 4.860 0.926 0.928 0.936 3.459 3.320 3.383 0.291 0.277 0.241
@@
Front-range HPE Results under Setting1. As shown in Table II, our SemiUHPE can surpass both the supervised methods [39] and baseline FisherMatch [36] with either backbone. Meanwhile, the smaller the ratio of labeled samples used (from 20% to 2%), the more significant the reduction in MAE errors. Moreover, with using 20% labeled 300W-LP, the performance of semi-supervised baseline and our SemiUHPE can always exceed the supervised method using all labeled 300W-LP. These results are also comparable to some SOTA supervised methods in Table III. We attribute this to the stronger robustness brought by partially integrated unsupervised training of unlabeled data. It also indicates that pure supervised learning may cause final models overfitting on the train-set, while SSL can improve the adaptability and generalization to a certain extent.
Front-range HPE Results under Setting2. As shown in Table III, with the supporting of SSL, our method based on ResNet50 can achieve a low MAE result 3.37, which is comparable to other supervised SOTA front-range HPE methods such as DAD-3DNet [16] with MAE 3.66, SynergyNet [10] with MAE 3.35 and DSFNet-f [61] with MAE 3.25. Please note that all of them are 3DMM-based which require dense face landmark labels and complex 3D face reconstruction pipelines. While, our SemiUHPE only requires reasonable excavation of unlabeled wild heads, which is more practical and scalable in real applications. When using a stronger backbone such as RepVGG as in 6DRepNet [20], our method can obtain a lower result with MAE 3.35, which further explains its generality and advantage. In particular, after adopting the superior backbone EfficientNetV2-S, we can further reduce the MAE errors into 3.31, which achieved the second best performance so far.
@@
Performance Improvement Analysis of SemiUHPE. To systematically understand the source of improvement after boosting FisherMatch with our proposed strategies, we further calculate and plot the head pose error gains on datasets AFLW2000 for front-range HPE under Setting2 (Fig. 10a) and DAD-3DHeads for full-range HPE under Setting3 (Fig. 10b). We split all the testing images into different groups according to their GT yaw angles and calculate the improved error degree within each group. We can find that the improvement by our SemiUHPE gets more obvious as the head pose becomes more challenging and has a larger yaw angle (e.g., $>90^{\circ}$). This capability benefits from unsupervised joint training using a large number of human heads collected in any orientation.
Results of SemiObjRot. The experiment results on Pascal3D+ dataset of generic object rotation regression are shown in Table VI. The results illustrate that, with using the effective teacher-student mutual learning framework as well as our proposed SSL strategies (including the dynamic entropy-based pseudo label filtering and two advanced strong augmentations), our SemiObjRot significantly outperforms various baselines and the state-of-the-art method FisherMatch under all different numbers of labeled images. This is strong evidence of the undisputed superiority and valuable versatility of our approach.
@@
4.4 Qualitative Comparison
To further explain the superiority of our SemiUHPE, we present the qualitative comparison with SOTA counterparts including SynergyNet [10], DAD-3DNet [16] and DSFNet-f [61]. As shown in Fig. 11 and Fig. 12, our method can obtain impressive HPE results on front yet severely occluded faces and in-the-wild challenging heads. Moreover, we present comparison on never-before-experienced yet challenging samples from the test-set of DAD-3DHeads. As shown in Fig. 13, although these images have high-definition, DAD-3DNet may make significant mistakes, leading to disordered results of head reconstruction. While, our method usually gives a satisfactory estimation. More convincing qualitative results can be found in our online project link https://hnuzhy.github.io/projects/SemiUHPE.

### v4

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
@@ 1 Introduction
In the era of deep supervised learning, many HPE methods [17, 18, 19, 20] need a large amount of labeled data to train. However, existing HPE datasets such as 300W-LP [11], BIWI [21] and DAD-3DHeads [16] have their incompatible limitations for real applications. They are either artificially collected or synthesized [11, 21, 22, 23] so that having huge domain gaps and scarce diversities compared to natural heads. Others are manually annotated by certified experts [16, 24] at a significant time and economic cost with a small scale. Some examples are shown in Fig. 2. In these labeled datasets, 300W-LP [11] covers yaw angles within $(-99^{\circ},99^{\circ})$ and BIWI [21] within $(-75^{\circ},75^{\circ})$, both focusing mainly on front-range visible faces. In contrast, DAD-3DHeads [16] officially reports 39% front, 52% side, and 9% atypical poses, providing broader coverage but still limited representation of backward and invisible heads. These statistics confirm that unconstrained, full-range head pose estimation remains largely underexplored.
Affected by this, although many supervised learning methods designed and trained based on these datasets have achieved excellent quantitative performance on the test set (such as AFLW2000 [11] and BIWI [21]), these models cannot be directly applied to the real world with complex scenarios. Some methods try to synthesize a large number of multi-view [25], extreme-view [26] or full-view [27] human heads to expand the training set, but these generated images still have hallucination defects and cannot be fully trusted. Different from most existing fully supervised methods, we turn to the semi-supervised learning (SSL) techniques [28, 29, 30, 31], and propose a semi-supervised unconstrained head pose estimation (SemiUHPE) method, that can leverage extensive easier obtainable yet unlabeled in-the-wild heads [32, 33, 34, 35] in addition to partially labeled data to improve performance and generalization. The overall framework of our method is illustrated in Fig. 3. It can not only avoid the laborious annotation of 3D head pose on 2D images which itself is ill-posed, but also greatly promote the estimation accuracy of challenging cases in real environments.
@@
The three above-mentioned strategies are complementary: cropping guarantees a stable and scale-invariant input, dynamic filtering ensures the reliability of semi-supervised training, and augmentations promote robustness against extreme pose and occlusion variations. Their synergy enables our framework to consistently improve accuracy and generalization in unconstrained HPE. In method section, we provide more details of them. We also present how these strategies can be utilized to solve the task generic object rotation regression similar to HPE and help improve another task 3D head reconstruction strongly related to HPE. Despite the simplicity, these seamless adaptations reveal valuable universality and scalability of SemiUHPE. In experiments, we demonstrate the effectiveness and versatility of our proposed strategies by extensive comparison and ablation studies. In addition, we give impressive qualitative estimation results on challenging wild heads, which is promising for downstream applications.
@@ 2.3 Semi-Supervised Rotation Regression
This is a less-studied field compared with other popular SSL tasks such as classification and detection. For the 6D pose estimation, Self6D [69, 70] establishes self-supervision for 6D pose by enforcing visual and geometric consistencies on top of unlabeled RGB-D images. NVSM [71] builds a category-level 3D cuboid mesh for estimating pose of rigid object in a synthesis-and-matching way. Recently, based on FixMatch [30], FisherMatch [36] firstly proposes the semi-supervised rotation regression for generic objects. It can automatically learn uncertainties along with predictions by introducing the matrix Fisher distribution [38, 39] to build a probabilistic model of rotation. The entropy of this distribution has been validated to be an efficient indicator of prediction for pseudo label filtering. The improved version [37] has integrated the rotation Laplace distribution [72] which is more robust to the disturbance of outliers and enforces much gradient to the low-error region. Essentially, HPE is subordinate to the rotation regression problem. MFDNet [56] also utilizes the matrix Fisher distribution to model head rotation uncertainty. Kuhnke et al. [73, 74] introduce the relative pose consistency into semi-supervised constrained HPE on dataset BIWI, while we explicitly target the more challenging unconstrained in-the-wild setting with novel cropping, dynamic filtering, and head-oriented augmentations. We follow [36, 56] and adopt this distribution to tackle the SemiUHPE task.
@@ 3 Method
Clarification of $\mathcal{D}^{u}_{id}$ and $\mathcal{D}^{u}_{ood}$. In this work, the terms $id$ and oodood are not intended to strictly follow the classical notions of covariate or distributional shift, but rather serve as task-specific operational definitions. Concretely, $\mathcal{D}^{u}_{id}$ denotes samples that, despite possibly having large head rotations, still provide sufficient discriminative cues (e.g.,, side-view heads with visible contours or head-top structure) such that the network prediction confidence remains relatively high. These heads are valuable for improving robustness although they are rarely covered by existing labeled datasets. By contrast, $\mathcal{D}^{u}_{ood}$ corresponds to extremely ambiguous cases with multiple sources of visual degradation (e.g., backward heads with heavy occlusion or blur), where even humans may struggle to infer orientation. Such cases are regarded as noisy and are progressively filtered by our dynamic entropy-based strategy.
@@ 3.1 Framework Overview
$\small p(\mathbf{R})=\mathcal{MF}(\mathbf{R};\mathbf{A})=\frac{1}{F(\mathbf{A})}\exp(\text{tr}(\mathbf{A}^{T}\mathbf{R}))$ (1)
@@
$\small\mathbf{R}=\mathbf{U}\left[\begin{array}[]{ccc}1&\qquad 0&\qquad 0\\ 0&\qquad 1&\qquad 0\\ 0&\qquad 0&\qquad\det(\mathbf{U}\mathbf{V})\end{array}\right]\mathbf{V}^{T}$ (2)
@@
$\small H(f)=\log{F_{f}}-\sum\nolimits^{4}_{i=1}\left({z_{f_{i}}\frac{1}{F_{f}}\frac{\partial{F_{f}}}{\partial{z_{f_{i}}}}}\right)$ (3)
@@ 3.3 Dynamic Entropy-based Filtering
$\small\mathcal{L}_{unsup}(\mathbf{x}^{u})=\mathds{1}_{(H(p_{tea})\leq\tau)}\mathcal{L}^{CE}(p_{tea},p_{stu})$ (4)
@@
$\small\tau_{k}=\mathsf{percentile}\langle H(\mathcal{MF}(\mathcal{N}_{tea}^{k}(\mathbf{x}^{u}_{i})))|_{i=1}^{N_{u}},\delta\rangle$ (5)
@@
$\small\mathcal{L^{\prime}}_{unsup}(\mathbf{x}^{u})=\mathds{1}_{(H(p^{k}_{tea})\leq\tau_{k})}\mathcal{L}^{CE}(p^{k}_{tea},p_{stu})$ (6)
where $p^{k}_{tea}\!=\!\mathcal{MF}(\mathcal{N}_{tea}^{k}(\mathbf{x}^{u}))$. Usually, with a given percentage $\delta$, the dynamic entropy threshold $\tau_{k}$ will decrease as the stage $k$ increases. And the optimal value of $\delta$ is inversely proportional to the images number of $\mathcal{D}^{u}_{ood}$ in $\mathcal{D}^{u}$. It should be emphasized that in our context, the distinction between $\mathcal{D}^{u}_{id}$ and $\mathcal{D}^{u}_{ood}$ reflects the difficulty and reliability of pose inference rather than classical covariate shift, which helps the dynamic threshold adaptively retain plausible but hard samples while filtering extremely noisy ones. Representative $\mathcal{D}^{u}_{id}$ and $\mathcal{D}^{u}_{ood}$ examples under our definition are visualized in Figs. 5, 11, 12, 14 to illustrate the filtering effect in practice.
@@ 3.4.2 Pose-Altering Rotation Consistency
$\overline{p^{i}_{stu}}=\mathcal{MF}(\mathcal{N}_{stu}(\overline{\mathbf{x}^{u}_{i}})),\qquad\overline{\mathbf{x}^{u}_{i}}=T_{\mathsf{CutOcc}}(T_{\mathsf{Rot}_{\theta}}(\mathbf{x}^{u}_{i}))$ (7)
$\widetilde{p^{i}_{tea}}=\mathcal{MF}(\mathcal{N}_{tea}(\widetilde{\mathbf{x}^{u}_{i}})),\qquad\widetilde{\mathbf{x}^{u}_{i}}=T_{\mathsf{weak}}(\mathbf{x}^{u}_{i})$
$\widehat{p^{i}_{tea}}=\mathbf{M_{\theta}}\widetilde{p^{i}_{tea}},\qquad\mathbf{M_{\theta}}=\small{\left[\begin{array}[]{ccc}\cos(\theta)&\sin(\theta)&0\\ -\sin(\theta)&\cos(\theta)&0\\ 0&0&1\end{array}\right]}$
@@
It is worth noting that Kuhnke et al. [73, 74] also employed rotation-based consistency within semi-supervised HPE, but their methods were designed for constrained benchmarks (e.g., BIWI), whereas our proposed $T_{\mathsf{CutOcc}}$ and $T_{\mathsf{Rot}_{\theta}}$ are specifically tailored for unconstrained in-the-wild heads with severe occlusion and atypical orientations.
@@ 3.5 Adaptation of SemiUHPE
Figure 8: Overall architecture of our Semi3DHead framework. It extends DAD-3DNet [16] by incorporating multiple input branches: (i) heatmaps and landmarks inherited from DAD-3DNet, (ii) a 3D Morphable Model (3DMM) branch using the FLAME model [81], and (iii) an additional $3\!\times\!3$ rotation matrix branch introduced in this work to directly regress head pose and filter out out-of-distribution samples.
@@
To filter unreliable pseudo-labeled heads, we introduce a new pose regression branch and design a pose-guided filtering strategy. As illustrated in Fig. 9 (left), one straightforward way is to measure the discrepancy between the predicted rotation matrix from the head-pose branch and the rotation matrix embedded in the 3DMM branch. This discrepancy can be quantified by the geodesic distance on the rotation group $\mathcal{SO}(3)$, which corresponds to the minimal angular difference between two rotations (approximated by the Frobenius norm $\|\mathbf{I}\!-\!\mathbf{R}_{1}\mathbf{R}_{2}^{T}\|_{F}$). A smaller geodesic distance indicates higher consistency and reliability. However, relying solely on matrix differences may not always provide a statistically robust measure of uncertainty. To address this, we further model the head pose prediction probabilistically by fitting it with a matrix Fisher distribution, which naturally captures the distribution of rotations on $\mathcal{SO}(3)$. In this setting, the distribution entropy serves as an indicator of prediction confidence: low entropy means the distribution is sharp and reliable, while high entropy corresponds to ambiguous or noisy samples. As shown in Fig. 9 (right), we use this entropy-based criterion to dynamically filter out high-uncertainty samples. There are two advantages to applying the matrix Fisher distribution. On one hand, its output is an arbitrary $3\!\times\!3$ matrix with homeomorphic parameters to $\mathbb{R}^{9}$ [39], which is easier to estimate than the 3D rotations in $\mathcal{SO}(3)$ topology [56]. On the other hand, we can measure the confidence of predicted distribution by calculating its entropy, and devise a self-reliant more robust filtering mechanism. Therefore, while different rotation representations can be mathematically transformed into each other, we adopt the $3\!\times\!3$ rotation matrix as the default representation. We will verify that the adapted SemiUHPE can benefit from both filtering strategies, with the latter being superior.
@@ 3.6 Training Protocol
where $\log(\cdot)$ means the negative log likelihood (NLL) of the mode predicted by $\mathcal{N}_{stu}$ in the distribution of ground-truth label $\mathbf{y}^{l}$. We save the best performed student model for cloning an identical teacher model for the next phase.
@@ 4.1 Datasets and Settings
Setting1: 300W-Self In this setting, we use partially annotated 300W-LP as the labeled set and the left part as the unlabeled set. The labeled ratio is selected from (2%, 5%, 10%, 20%). The Phase1 and Phase2 has 20K and 40K iterations, respectively. The number of threshold updating stages $K$ in Phase2 is 4. Batch size for the labeled set $B_{l}$ and unlabeled set $B_{u}$ is 32 and 128, respectively. The remained percentage $\delta$ of unlabeled data is 0.95. The test-set is AFLW2000. For a fair comparison, we keep samples with Euler angles within $\pm 90^{\circ}$ following the previous front-range HPE methods [18, 19, 20]. The evaluation metric is Mean Absolute Error (MAE) of Euler angles. We thus convert predictions into Euler angles for comparing.
@@
Compared Methods: In order to make a more comprehensive comparison, apart from our proposed SemiUHPE, we also implemented the fully supervised version (Sup.) using only labeled data but adopting an advanced matrix Fisher representation [39], and the semi-supervised baseline FisherMatch [36] (Base.) following our settings. For the network backbone, it is selected from three purely ConvNets-based candidates including ResNet50 [87], RepVGG [88] and EfficientNetV2-S [89]. These backbones have similar network parameters and are widely used in other HPE methods [18, 19, 20]. Besides, we also included two additional baselines to strengthen the comparisons: (1) FisherMatch+ [37] (Base.+), where we replace the matrix Fisher distribution with a rotation Laplace distribution; (2) Sample Adaptive Augmentation [90] (SAA), an SSL method from image classification that introduces stronger adaptive augmentations. Both Base.+ and SAA were re-implemented on top of the best-performing settings of FisherMatch for SemiUHPE. For all experiments, we implemented them using PyTorch 1.13.1 on one single RTX 3090 (24 GB) or A800 (80 GB) GPU. The Adam optimizer is used. The learning rate in Phase1 is 1e-4, and reduced to 1e-5 in Phase2. During training and testing, all head images are padded and resized into shape $224\times 224$ as inputs.
@@ 4.2 Implementation Details
$T_{\mathsf{CutOcc}}$
CutOut: 3 Holes
CutMix: 3 Holes
1.0
@@ 4.3 Quantitative Comparison
Below, we report results under three settings as well as improvement analysis of our SemiUHPE. If there is no special reminder, all baselines (Sup., Base., Base.+ and SAA) are implemented with our proposed aspect-ratio invariant cropping. The last is the results of SemiObjRot and Semi3DHead.
TABLE II: Euler angles errors on AFLW2000. Models are trained on 300W-LP with different ratios of label. The best result is in red color. SL and SSL are fully supervised learning and semi-supervised learning, respectively.
  Type Method Backbone 2% 5% 10% 20% All
  SL Sup. ResNet50 4.347 3.987 3.831 3.619 3.578
@@
TABLE III: Euler angles errors on AFLW2000. Models are all trained on 300W-LP. Extra means additional annotations (e.g., Landmarks (LMs) or 3DMM). The marker $\dagger$} indicates using additional labeled training data. The best and second-best result is in red and blue color, respectively.
  Type Method Reference Extra Pitch Yaw Roll MAE
  SL 3DDFA [11] CVPR’16 3DMM 5.98 4.33 4.30 4.87
@@
SSL Base. (ResNet50) CVPR’22 No 4.61 2.99 3.00 3.53
@@
[…]
@@
[…]
@@
  SL DAD-3DNet [16] 7.861 4.766 4.025 2.467 0.919 0.929 0.938 0.954 3.957 3.468 3.247 2.982 0.327 0.270 0.233 0.149
SSL Mean-Teacher [28] 5.599 4.969 4.860 0.926 0.928 0.936 3.459 3.320 3.383 0.291 0.277 0.241
@@
Front-range HPE Results under Setting1. As shown in Table II, our SemiUHPE can surpass both the supervised methods [39] and baseline FisherMatch [36] with either backbone. When using ResNet50, our SemiUHPE is always the best out of other three SSL methods (Base., Base.+ and SSA). Meanwhile, the smaller the ratio of labeled samples used (from 20% to 2%), the more significant the reduction in MAE errors. Moreover, with using 20% labeled 300W-LP, the performance of semi-supervised baseline and our SemiUHPE can always exceed the supervised method using all labeled 300W-LP. These results are also comparable to some SOTA supervised methods in Table III. We attribute this to the stronger robustness brought by partially integrated unsupervised training of unlabeled data. It also indicates that pure supervised learning may cause final models overfitting on the train-set, while SSL can improve the adaptability and generalization to a certain extent.
Front-range HPE Results under Setting2. As shown in Table III, with the supporting of SSL, our method based on ResNet50 can achieve a low MAE result 3.37, which is comparable to other supervised SOTA front-range HPE methods such as DAD-3DNet [16] with MAE 3.66, SynergyNet [10] with MAE 3.35 and DSFNet-f [61] with MAE 3.25. Please note that all of them are 3DMM-based which require dense face landmark labels and complex 3D face reconstruction pipelines. While, our SemiUHPE only requires reasonable excavation of unlabeled wild heads, which is more practical and scalable in real applications. Comparing to other three SSL baselines (Base., Base.+ and SSA), our method still outperforms them under this full-range setting. When using a stronger backbone such as RepVGG as in 6DRepNet [20], our method can obtain a lower result with MAE 3.35, which further explains its generality and advantage. In particular, after adopting the superior backbone EfficientNetV2-S, we can further reduce the MAE errors into 3.31, which achieved the second best performance so far.
@@
Performance Improvement Analysis of SemiUHPE. To systematically understand the source of improvement after boosting FisherMatch with our proposed strategies, we further calculate and plot the head pose error gains on datasets AFLW2000 for front-range HPE under Setting2 (Fig. 10a) and DAD-3DHeads for full-range HPE under Setting3 (Fig. 10b). We split all the testing images into different groups according to their GT yaw angles and calculate the improved error degree within each group. We can find that the improvement by our SemiUHPE gets more obvious as the head pose becomes more challenging and has a larger yaw angle (e.g., $90^{\circ}$). This capability benefits from unsupervised joint training using a large number of human heads collected in any orientation.
Results of SemiObjRot. The experiment results on Pascal3D+ dataset of generic object rotation regression are shown in Table VI. The results illustrate that, with using the effective teacher-student mutual learning framework as well as our proposed SSL strategies (including the dynamic entropy-based pseudo label filtering and two advanced strong augmentations), our SemiObjRot significantly outperforms various baselines and three state-of-the-art methods (including FisherMatch, FisherMatch+, and the adapted SAA) under all different numbers of labeled images. This is strong evidence of the undisputed superiority and valuable versatility of our approach.
@@
4.4 Qualitative Comparison
To further explain the superiority of our SemiUHPE, we present the qualitative comparison with SOTA counterparts including SynergyNet [10], DAD-3DNet [16] and DSFNet-f [61]. As shown in Fig. 11 and Fig. 12, our method can obtain impressive HPE results on front yet severely occluded faces and in-the-wild challenging heads. Moreover, we present comparison on never-before-experienced yet challenging samples from the test-set of DAD-3DHeads. As shown in Fig. 13, although these images have high-definition, DAD-3DNet may make significant mistakes, leading to disordered results of head reconstruction. While, our method usually gives a satisfactory estimation. More convincing qualitative results can be found in our project link.